3x+6y=15
-2x+3y=-3

Answer by mananth(16946)

3 x + 6 y = 15 .............1
-2 x + 3 y = -3 .............2
Eliminate y
multiply (1)by 1
Multiply (2) by -2
3 x + 6 y = 15
4 x + -6 y = 6
Add the two equations
7 x = 21
/ 7
x = 3
plug value of x in (1)
3 x + 6 y = 15
9 + 6 y = 15
6 y = 15 -9
6 y = 6
y = 1